OpenStudy (anonymous):

how do u find the area and circumfrance of a circle with the diameter of 18 inches. Express your answer in pie?

OpenStudy (anonymous):

A=\[\pi\]r(squared)(times2) C=2\[\pi\]r r stands for radius so the radius would be half of the diameter 18=9 and now you just solve it

OpenStudy (amistre64):

C = (d)pi = 18pi A = C/2 = 18pi/2 = 9pi

OpenStudy (amistre64):

make that A = Cr/2 = 9(9)pi/2 = 81pi/2

OpenStudy (amistre64):

forget it...lol...i gotta take a nap :)